OpenStudy (anonymous):

Police use the formula: v = sqrt(20 L) to estimate the speed of a car, v, in miles per hour, based on the length, L, in feet, of its skid marks when suddenly braking on a dry, asphalt road. At the scene of an accident, a police officer measures a car's skid marks to be 188 feet long. Approximately how fast was the car traveling? Round your answer to the nearest tenth (one decimal place) of a unit.

OpenStudy (sweetburger):

Police use the formula gg

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (triciaal):

v = sq rt (20 * 188) v = sq rt (4*5*4* 47) = 4 rt5*rt47 estimated v = 4*2*7= 56 actual = sq rt 3760 then round to 1 decimal

OpenStudy (anonymous):

That was it. Can you explain how you got to the answer?

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

They wanted the actual answer with 1 decimal place

OpenStudy (triciaal):

i did look at the given formula replace l with 188 and solve

OpenStudy (triciaal):

did you get it or do you need more help?

OpenStudy (anonymous):

I got it. thanks :)